OpenStudy (anonymous):

For a quadratic function, f(1)=5, f(2)=-7, f(3)=-11 find f(-2) I'm studying for my exam and do not know how to do this. Please help me and tell me step by step. Thank you!

OpenStudy (asnaseer):

do you know the general form for a quadratic equation?

OpenStudy (anonymous):

isn't it ax^2+bx+c

OpenStudy (asnaseer):

thats right. so you know your function is of the form:\[f(x)=ax^2+bx+c\]

OpenStudy (asnaseer):

a, b and c are three unknowns here

OpenStudy (asnaseer):

and you have been given three values for f(x)

OpenStudy (asnaseer):

you should be able to use these values to find a, b and c

OpenStudy (anonymous):

so what i did was a+b+c=5 4a+2b+c=-7 9a+3b+c=-11 but Im not sure what to do next to find f(-2)

OpenStudy (asnaseer):

use the 3 equations to solve for a, b and c. then use these values of a, b and c to find f(-2)

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

would a matrices help?

OpenStudy (asnaseer):

you could use matrices, but these equations are fairly simple. e.g.: a+b+c=5 (1) 4a+2b+c=-7 (2) 9a+3b+c=-11 (3) 3a+b=-12 (2) - (1) 5a+b=-4 (3) - (2)

OpenStudy (asnaseer):

then eliminate b from the 2 new equations and solve for a then find b then find c

OpenStudy (anonymous):

so i got f(-2)=-535…but i don't think that is right

OpenStudy (anonymous):

i got a=4, b=-24, c=25

OpenStudy (asnaseer):

those are correct

OpenStudy (asnaseer):

so you now have:\[f(x)=4x^2-24x+25\]

OpenStudy (asnaseer):

how did you calculate f(-2)?

OpenStudy (anonymous):

so i now got 89 i think i did something wrong on my calculator

OpenStudy (asnaseer):

yup - that's the correct answer - well done! :)

OpenStudy (anonymous):

Thank you so much for your help!!

OpenStudy (asnaseer):

yw :)